i need to find the number of square yards of sod needed to cover the circular putting green shown
(hint:9sq ft= 1 sq yd)
the circle has a radius of 15 ft

Answer by checkley77(12844)

15ft=5yds.
a=pir^2
a=3.14*5^2
a=3.14*25
a= 78.5 yds^2.